OpenStudy (anonymous):

For triangle ABC with c=345, a=534, and C=25.4 degree, there are two possible values for angle A. What are they?

OpenStudy (johnweldon1993):

 We can use the law of sines there \[\large \frac{sin(25.4)}{345} = \frac{A}{534}\] And solve for A

OpenStudy (anonymous):

i got it but it has 2 different values man

OpenStudy (anonymous):

one i got is 44.43 degree

OpenStudy (anonymous):

any clue guys?

OpenStudy (anonymous):

How did you even get 44.3? I got 41.6 degrees.

OpenStudy (anonymous):

The other answer could therefore be (180-41.6)= 138.4 degrees.

OpenStudy (anonymous):

lol...my bad...the correct answer is 41.6

OpenStudy (johnweldon1993):

Hmm...I'm not entirely sure on that It states that angle A can have 2 answers...since angle A will always oppose side 'a' it will always be in the same ratio plugged into the law of sines...it can never be what angle B will be which would be the 138.4 degrees Frankly I'm not sure how 1 angle can have 2 different values if side lengths are given as they are

OpenStudy (anonymous):

this is the point that confused me, but this is the question i got in book...so....

OpenStudy (wolf1728):

Guess I should post something here. When you are given two sides and a NON-INCLUDED angle, two triangles can sometimes result. (see graphic). Notice that the 3 items of given information - (25.4° angle and sides of 534 and 345) remain the same in both triangles. However, angles A and B and side b do change. Just thought I'd show how you can get two triangles when given 2 sides and a non-included angle, More information and a calculator are located here: http://www.1728.org/trigssa.htm